INVOLUTIONS ON MODULI SPACES AND REFINEMENTS OF THE VERLINDE FORMULA

JØRGEN ELLEGAARD ANDERSEN AND GREGOR MASBAUM

Abstract. The moduli space M of semi-stable rank 2 bundles with trivial determinant over a complex curve Σ carries involutions naturally associated to 2-torsion points on the Jacobian of the curve. For every lift of a 2-torsion point to a 4-torsion point, we define a lift of the involution to the determinant line bundle L. We obtain an explicit presentation of the group generated by these lifts in terms of the order 4 Weil pairing. This is related to the triple intersections of the components of the fixed point sets in M, which we also determine completely using the order 4 Weil pairing. The lifted involutions act on the spaces of holomorphic sections of powers of L, whose dimensions are given by the Verlinde formula. We compute the characters of these vector spaces as representations of the group generated by our lifts, and we obtain an explicit isomorphism (as group representations) with the combinatorial-topological TQFT-vector spaces of Blanchet et al. As an application, we describe a ‘brick decomposition’, with explicit dimension formulas, of the Verlinde vector spaces. We also obtain similar results in the twisted (i.e., degree one) case.